A shopkeeper buys a crate of eggs at $1.50 per dozen. he buys another crate, containing 3 dozen more than the first crate, at $2.00 per dozen. He sells them all for $2.50 a dozen and makes $15 profit. How many dozens were there in each of the crates?
Answer by mananth(16949)

A shopkeeper buys a crate of eggs at $1.50 per dozen. he buys another crate, containing 3 dozen more than the first crate, at $2.00 per dozen. He sells them all for $2.50 a dozen and makes $15 profit. How many dozens were there in each of the crates?
let the eggs be x dozens
cost of eggs = 1.5x
Crate II.
x+3 dozens
cost = 2*(x+3)
Cost of eggs = 1.5x+2(x+3)
= 1.5x+2x+6
=3.5x+6
Number of eggs = x+(x+3)
sold for 2.50
Sales = 2.50(2x+3)
=5x+7.5
Sales - cost = 15
2.50(2x+3)-1.5x+2(x+3)=15
5x+7.50-3.5x-6=15
1.5x=13.5
x=9
Crate I--- 9 dozens
Crate II---- 12 dozens